OpenStudy (anonymous):

prove that: sin^2A - cos^2A.cos2B = sin^2B - cos^2B.cos2A

OpenStudy (aravindg):

Okay start by expressing cos 2B as cos^2B-sin^2B Sin^2A-cos^2A.(cos^2B-sin^2B) =sin^2A-cos^2A.cos^2B+cos^2A.sin^22B Now add and subtract cos^2B.sin^2A here After that cos^2B as common and you get cos^2B.cos2A The rest is easy: =sin^2A-cos^2B.cos 2A-cos^2Bsin^2A+cos^2Asin^2B =sin^2Asin^2B+cos^2Asin^2B-cos^2B.cos2A =Sin^2B-cos^2B.cos 2A=RHS Hence proved

OpenStudy (aravindg):

Are you able to follow?

OpenStudy (anonymous):

=sin^2A-cos^2B.cos 2A-cos^2Bsin^2A+cos^2Asin^2B how did this come???

OpenStudy (aravindg):

By doing the process I said. I didn't show it as I thought you can do it yourself.

OpenStudy (anonymous):

actually, i didn't understand,, add and subtract cos^2B.sin^2A means??

OpenStudy (aravindg):

=sin^2A-cos^2A.cos^2B+cos^2A.sin^2B+(cos^2B.sin^2A )-(cos^2B.sin^2A ) Do you get it now?

OpenStudy (anonymous):

oh oh.. i thought of something else :p

OpenStudy (aravindg):

lol ok Can you do the next step and get to =sin^2A-cos^2B.cos 2A-cos^2Bsin^2A+cos^2Asin^2B ??

OpenStudy (anonymous):

yeah.. thanks :)

OpenStudy (aravindg):

Great! Yw :)

OpenStudy (anonymous):

will u solve my another question too? i have already posted it!!

OpenStudy (aravindg):

Sure but after 20 minutes maybe...Time for dinner!

OpenStudy (anonymous):

ok...no problem..